# Edge-Spreading Raptor-Like LDPC Codes for 6G Wireless Systems

### v1

@@ III Edge-Spreading Raptor-Like LDPC Codes
Figure 2: Uncoupled protograph and protomatrix for an ESRL code with a HRC, an IRC, and an SSC. Upper right corner are empty to ensure rate compatibility.
@@ III-A Structure of ESRL Codes
Fig. 2 illustrates the structures of the uncoupled protograph $\mathcal{G}$ and uncoupled protomatrix $\mathbf{B}$ of an ESRL code, which consists of three components: a highest-rate code (HRC), an incremental redundancy code (IRC), and a supplementary-spreading code (SSC). Unlike conventional protographs of PBRL codes (such as the ones used in 5G-NR) that feature a rightmost identity matrix [17], an SSC matrix is used to maximize the possible locations for edge placement. Let (1) denote $\mathbf{B}$ of an ESRL code
@@
Previous approaches to designing rate-compatible SC-LDPC codes [31, 32, 33, 34, 35] typically design a good uncoupled protomatrix first, followed by coupling. For instance, [33, 34, 35] attempted to generate rate-compatible SC-LDPC codes by directly applying edge spreading to a PBRL protograph. However, as shown in Fig. 3, decoding of SC-LDPC codes now occurs on the coupled $\mathbf{B}_{\text{SC}}$ rather than the uncoupled $\mathbf{B}$ like in LDPC block codes. This highlights a fundamental difference of our ESRL codes compared to [31, 32, 33, 34, 35], i.e., our target is to design an uncoupled protomatrix $\mathbf{B}$ that ultimately results in a good coupled protomatrix $\mathbf{B}_{\text{SC}}$ after edge spreading, regardless of whether $\mathbf{B}$ by itself is optimal. Given the Raptor-like structure inherent to ESRL codes, establishing a good Raptor-like structure after edge spreading is clearly more important than good performance before coupling. Define a partially coupled chain as
@@
Given the convolutional structure of SC-LDPC codes, decoding can be further decomposed into multiple identical engines (shown in Fig. 3) working on different $\mathbf{U}$ blocks at various spatial positions in parallel (also called multiple-engine decoding in [12]). Intuitively, decoding on $\mathbf{U}$ largely determines the overall performance of SC-LDPC decoding. However, a conventional PBRL structure restricts the design space to find the optimal $\mathbf{B}_{\text{SC}}$, since its rightmost SSC is only an identity matrix and must be fully allocated to $\mathbf{B}_{0}$ for encoding [37]. This reduces the possible locations for edge placement and leaves the right parts of $\mathbf{B}_{\omega},\mathbf{B}_{\omega-1},\dots,\mathbf{B}_{1}$ in (2) empty. The proportion of empty parts increases to $67.6\%$ across medium to low rates, leading to performance loss [33, 34, 35]. Hence, SC-LDPC codes need a new structure for the uncoupled protomatrix to enhance message propagation on $\mathbf{U}$. Given that [14, 15, 16] have indicated that all rate-compatible LDPC codes must adhere to a lower triangular constraint on the right side, either uncoupled or coupled, the ESRL code also follows this constraint to preserve the upper right corner of (1) empty and keep the SSC lower triangular. Note that despite the motivation of ESRL codes to design the optimal $\mathbf{B}_{\text{SC}}$ after edge spreading, the ESRL code still retains the fundamental HRC and IRC parts of traditional PBRL codes. This means an ESRL decoder can be naturally compatible with PBRL codes of appropriate sizes such as 5G-NR LDPC codes.
The ESRL structure also resembles Raptor codes [21], but with key differences. First, while the HRC is identical to the precode in a Raptor code, the ESRL code is systematic, with the codeword of each batch represented as $\bm{c}_{[i]}=[\bm{u}_{[i]},\bm{r}_{[i]}],0\leq i<L$. This design allows explicit transmission of the HRC VNs in all batches and supports decoding at the highest rate, which is infeasible in a Raptor code. Second, similar to the LT component in Raptor codes, XOR operations in the ESRL code generate the VNs of the IRC and the SSC, producing additional parity bits row-wise. However, these connections in the ESRL code are pre-defined rather than arbitrary. Third, compared to two-phase BP decoding of Raptor codes or joint BP decoding on PBRL codes, the ESRL code can support a more flexible windowed BP decoding to slide over the coupled protograph $\mathcal{G}_{\text{SC}}$ in a low-complexity manner [38]. More details on decoding are discussed in Section V-B.
@@ III-B Properties of ESRL Codes
In this section, we present some important properties of ESRL codes. First, the ESRL code is a time-invariant SC-LDPC code with no parallel edges, i.e., each entry of $\mathbf{B}$ is binary, and no non-zero entries in $\mathbf{B}_{i},\forall i\in[0,\omega]$ occupy the same position. The time-invariant nature and the absence of parallel edges ensure that the ESRL code can be decomposed into multiple independent identical engines that decode on different $\mathbf{U}$ blocks in parallel, without causing memory access conflicts (assuming the decoding schedule is consistent at each engine). This results in ESRL codes that allow for significant decoding parallelism to potentially achieve the high throughput demand of next-generation channel coding, as demonstrated in Section V-B and Section VI. It is noteworthy that although parallel edges can reduce the decoding threshold [17, 39] and can be easily assigned to $\omega$ sub-matrices in SC-LDPC codes without the need for two-step lifting [39], this performance gain sacrifices parallelism in an ESRL decoder. The absence of parallel edges, as mentioned in [29], can also simplify the theoretical tools described in Section IV to optimize ESRL codes.
The HRC protograph is the protograph design of a good code at a given rate, which has been well studied in [39]. Following the example of [39], we set the minimum column degree to at least $3$ to ensure linear minimum distance growth with increasing blocklength. Besides, a punctured VN that is fully connected to all CNs in the HRC is placed in the first column to reduce the decoding threshold. The IRC component aims to provide reliability back to the VNs of the original HRC. In a complete ESRL matrix with the lowest rate, the CNs of the IRC must eventually connect back to all VNs of the HRC. Following [40], each row of the IRC has at least two edges to retain linear minimum distance growth with increasing blocklength. Besides, the punctured VN of the precode connects to all CNs in the IRC. Unlike the use of parallel edges in [17], the HRC and the IRC of ESRL codes do not involve any parallel edges.
The SSC protograph is to enhance message propagation on $\mathbf{U}$, especially for medium to low rates. In the SSC, all diagonal elements are set to ‘$1$’s to facilitate encoding. Note that the SSC is connected to and primarily improves the reliability of VNs associated with redundancy, which are generally less critical than the VNs in the HRC, thus the weight of the SSC should be kept lower than that of the IRC. While the rate reduces, we increase the SSC weight to avoid the right side being too sparse, but by experimental trial, this proportion is at most $30\%$ of the IRC weight.
@@ III-C Code Profile of ESRL Codes
The uncoupled protomatrix $\mathbf{B}$ outlines the parent structure of ESRL codes, as explained in Section III-A. To convert $\mathbf{B}$ to $\mathbf{B}_{\text{SC}}$, following [30, 29], we introduce an edge-spreading matrix $\mathbf{T}$, instead of explicitly recording $\omega+1$ sub-matrices as in [23, 26]. The matrix $\mathbf{T}$ of size $m\times n$ determines how the edges of the uncoupled protograph $\mathcal{G}$ are spread to the coupled protograph $\mathcal{G}_{\text{SC}}$. Specifically, the elements of $\mathbf{T}$ indicate to which sub-matrices the edges of $\mathbf{B}$ are assigned. We set $\mathbf{T}_{c,v}=-1$ if $\mathbf{B}_{c,v}=0$. Otherwise, $\mathbf{T}_{c,v}=i$, $i\in[0,\omega]$, implies that the entry at row $c$ and column $v$ of $\mathbf{B}_{c,v}$ is assigned to the corresponding position in $\mathbf{B}_{i}$. Let $\delta_{i}(\mathbf{T}_{c,v})$ be a function that returns a boolean value of ‘$1$’ if $\mathbf{T}_{c,v}$ equals $i$, and ‘00’ otherwise. As a result, we can generate $\omega+1$ sub-matrices $\mathbf{B}_{0},\mathbf{B}_{1},\dots,\mathbf{B}_{\omega}$ explicitly as
@@
Furthermore, the termination of finite-length SC-LDPC codes results in an unavoidable rate loss at the boundary. To mitigate this issue, one can either append a redundancy matrix $\mathbf{Q}$ at the end [41, 37] or reduce the information length of the last block [42] to satisfy encoding. ESRL codes adopt the former approach, as it has a lower rate loss. Previous works generally use a simple identity matrix for $\mathbf{Q}$ [41, 37], but this leads to serious error floors at medium to low rates. In contrast, we adopt a tri-diagonal matrix for $\mathbf{Q}$ with a lower triangular structure to ensure that most column degrees are either $2$ or $3$. For rate compatibility, the matrix $\mathbf{Q}$ must remain full-rank even after pruning any $j$ rows and $j$ columns from the end. Based on the three matrices $\mathbf{B}$, $\mathbf{T}$, and $\mathbf{Q}$, we explicitly define the coupled protomatrix $\mathbf{B}_{\text{SC}}$ of the ESRL code as follows
@@ IV-A Unified Graph for ESRL Codes
Inspired by the aforementioned code profile, we propose a novel graph representation for ESRL codes, which we call the unified graph $\mathcal{G}_{\text{U}}$. This graph contains all essential code information and enables a comprehensive illustration of all possible graph structures within an SC-LDPC protograph. Unlike traditional approaches, a single unified graph $\mathcal{G}_{\text{U}}$ can integrate the relationships and dependencies among the matrices $\mathbf{B}$ and $\mathbf{T}$ in ESRL codes, [^1] Given that the tail matrix $\mathbf{Q}$ has a fixed structure and is not involved in edge spreading, we ignore the presence of $\mathbf{Q}$ to simplify the optimization on $\mathcal{G}_{\text{U}}$. Note that $\mathcal{G}_{\text{U}}$ is a protograph concept. thereby offering a more global view of the code construction. The construction of $\mathcal{G}_{\text{U}}$ is built on the bipartite graph derived from $\mathbf{B}$, where edges are annotated by different line types based on the corresponding elements of $\mathbf{T}$. For example, Fig. 4 depicts the unified graph $\mathcal{G}_{\text{U}}$ derived from $\mathbf{B}$ and $\mathbf{T}$ with $\omega=1$, where dashed and solid lines distinguish whether edge $e$ is assigned to $\mathbf{B}_{0}$ or $\mathbf{B}_{1}$, respectively. In this section, we present a message-passing algorithm on $\mathcal{G}_{\text{U}}$ to count and eliminate short cycles of $\mathcal{G}_{\text{SC}}$ efficiently, which inspires a good spreading strategy on $\mathcal{G}_{\text{U}}$ to enhance the performance of ESRL codes.
@@ IV-A1 Cycle Counting over the Unified Graph
For rigour, we only consider simple cycles in this paper, i.e., tailless backtrackless closed (TBC) walks in a graph. After edge spreading from $\mathcal{G}$ to $\mathcal{G}_{\text{SC}}$, we can get Remark 1.
Remark 1.
The number of $l$-cycles in $\mathcal{G}$ is the upper bound of the number of $l$-cycles in $\mathcal{G}_{\text{SC}}$ if repeated cycles in $L$ replicas are counted only once.
@@
Subsequently, we introduce a message-passing algorithm over $\mathcal{G}_{\text{U}}$ to efficiently count cycles for ESRL codes. In general, a message-passing algorithm computes messages at the nodes and propagates them along the edges of a graph, where the calculation of the message sent along an edge $e$ excludes the message previously received along $e$ (refers to this property as extrinsic[7]. Let $\gamma$ denote a basic message, and an example of an extrinsic message-passing process is illustrated in Fig. 4. Hence, the messages over the edges of $\mathcal{G}_{\text{U}}$ are all monomials of $\gamma$ and the message-passing operation is multiplication. Besides, we need to record the extra accumulated path value $\phi$ of each $\gamma$ over $\mathcal{G}_{\text{U}}$, due to using edge spreading. Let $t$ denote the index of discrete time, and we call each $t$ as one step, corresponding to either a VN-to-CN or CN-to-VN phase, where $t$ always starts from 00 of the VN-to-CN phase. We denote $\mathcal{M}_{v,c}^{(t)}$ as the VN-to-CN message along $e_{v}^{c}$ and $\mathcal{M}_{c,v}^{(t)}$ as the CN-to-VN message along $e_{c}^{v}$. The update rules are defined in (9) and (10), and the $\phi$ of each message is also tracked dynamically to verify whether the cycle remains or not in $\mathcal{G}_{\text{SC}}$ based on Remark 1.
Herein, we present how to use $\mathcal{G}_{\text{U}}$ count cycles that include a specific VN. For the $\bar{v}$-th VN, the initial message $\mathcal{M}_{\bar{v},c}^{(0)}$ at $t=0$ along the edge $e_{\bar{v}}^{c}$ is denoted by $\gamma_{c}^{\phi_{c}}$, where $\phi_{c}$ is initialized as $\mathbf{T}_{c,\bar{v}}$. We can summarize the VN-to-CN message-passing rules as
@@
Notably, the general message format $\mathcal{M}$ that propagates along the edges in $\mathcal{G}_{\text{U}}$ is the product of multiple $\gamma$ variables, each labeled by individual path value $\phi$, where both the number and type of $\gamma$ are dynamic. At the $t\!=\!l-1$-th step, we will obtain the number of $l$-cycles that include the $\bar{v}$-th VN as
$Q_{\bar{v}}^{l}=f\left(p_{\bar{v}}^{l})\right/2,$ (13)
where the division by $2$ arises from the directed edges, as each cycle is counted twice. The function $f\left(p_{\bar{v}}^{l}\right)$ computes the number of all $\gamma$ variables with $\phi=0$ in the polynomial $p_{\bar{v}}^{l}=\sum_{c\in\mathbb{C}_{\bar{v}}}\mathcal{M}_{c,\bar{v}}^{(l-1)}$, i.e., only sequences with $\phi=0$ will remain as cycles in $\mathcal{G}_{\text{SC}}$ after edge spreading. Note that due to the extrinsic property mentioned before, the component formed by the initial message $\gamma_{c}$ sent along the edge $e_{\bar{v}}^{c}$ should be excluded from $\mathcal{M}_{c,\bar{v}}^{(l-1)}$ in (13[43]. Consequently, we repeat the above procedures for all VNs, $v\in\mathbb{V}$. The total number of $l$-cycles in the entire $\mathcal{G}_{\text{SC}}$ can be obtained as (14), where the accumulated result in (14) needs to be divided by $l$ as every $l$-cycle is counted $l$-times. Define the shortest cycle length in a graph as the girth $g$. The above approach can efficiently count $l$-cycles when $l\in\{4,6,\dots,2g-2\}$.
$Q^{l}=\left(\sum\limits_{v\in\mathbb{V}}Q_{v}^{l}\right)/\left(l/2\right).$ (14)
@@
    At $t=0$, $v_{4}$ sends the initial messages $\gamma_{1}^{1}$$\gamma_{2}^{0}$, and $\gamma_{3}^{0}$ along the adjacent edges $e_{v_{4}}^{c_{1}}$$e_{v_{4}}^{c_{2}}$, and $e_{v_{4}}^{c_{3}}$, respectively, where the initial path values are according to the corresponding elements of $\mathbf{T}$. All the messages sent by $v_{0}$, $v_{1}$, $v_{2}$, and $v_{3}$ along their edges are equal to $1$ and omitted. At $t=1$, all valid messages are updated based on (10) and (12) for the CN-to-VN phase.

### v2

@@ III-A Structure of ESRL Codes
Fig. 2 illustrates the structures of the uncoupled protograph $\mathcal{G}$ and uncoupled protomatrix $\mathbf{B}$ of an ESRL code, which consists of three components: an HRC $\mathbf{B}_{\text{HRC}}$, an IRC $\mathbf{B}_{\text{IRC}}$, and a supplementary-spreading code (SSC) $\mathbf{B}_{\text{SSC}}$. Unlike conventional protographs of PBRL codes [18], such as those used in 5G-NR, which include a rightmost identity matrix, an SSC matrix is used to maximize the possible locations for edge placement. Let (2) denote $\mathbf{B}$ of an ESRL code
@@
Previous approaches to designing rate-compatible SC-LDPC codes [40, 41, 42, 43, 44] typically design a good uncoupled protomatrix first, followed by coupling. For instance, [42, 43, 44] attempted to generate rate-compatible SC-LDPC codes by directly applying edge spreading to a PBRL protograph. However, as shown in Fig. 3, decoding of SC-LDPC codes now occurs on the coupled $\mathbf{B}_{\text{SC}}$ rather than the uncoupled $\mathbf{B}$ like in LDPC block codes. This highlights a fundamental difference of ESRL codes compared to [40, 41, 42, 43, 44], i.e., our target is to design an uncoupled protomatrix $\mathbf{B}$ that ultimately results in a good coupled protomatrix $\mathbf{B}_{\text{SC}}$ after edge spreading, regardless of whether $\mathbf{B}$ by itself is optimal. Given the Raptor-like structure inherent to ESRL codes, establishing a good Raptor-like structure after edge spreading is clearly more important than good performance before coupling. With the convolutional structure of SC-LDPC codes, decoding can be further decomposed into multiple identical engines (shown in Fig. 3) working on different $\mathbf{U}$ blocks at various spatial positions in parallel (also called multiple-engine decoding in [13]), where a partially coupled chain $\mathbf{U}$ is defined as
@@
Intuitively, decoding on $\mathbf{U}$ largely determines the overall performance of SC-LDPC decoding. However, a conventional PBRL structure restricts the design space to find the optimal $\mathbf{B}_{\text{SC}}$, since its rightmost SSC is only an identity matrix that must be fully allocated to $\mathbf{B}_{0}$ for encoding [46]. This reduces the possible locations for edge placement and leaves the lower triangular parts of $\mathbf{B}_{\omega},\mathbf{B}_{\omega-1},\dots,\mathbf{B}_{1}$ in (3) empty. The proportion of empty parts increases to $67.6\%$ across medium to low rates, leading to performance loss [42, 43, 44]. Hence, SC-LDPC codes need a new structure for the uncoupled protomatrix to enhance message propagation on $\mathbf{U}$. Given that [15, 16, 17] have indicated that rate-compatible LDPC codes must adhere to a lower triangular constraint on the right side, either uncoupled or coupled, the ESRL code also follows this constraint to preserve the upper right corner of (2) empty and keep the SSC lower triangular. Note that despite the motivation of ESRL codes to design the optimal $\mathbf{B}_{\text{SC}}$ after edge spreading, the ESRL code still retains the fundamental HRC and IRC parts of traditional PBRL codes. This means an ESRL decoder can be naturally compatible with PBRL codes of appropriate sizes such as 5G-NR LDPC codes.
The ESRL structure also resembles Raptor codes [24], but with key differences. First, while the HRC is identical to the precode in a Raptor code, the ESRL code is systematic, with the codeword of each batch represented as $\bm{c}_{[i]}=[\bm{u}_{[i]},\bm{r}_{[i]}],0\leq i<L$. This design allows explicit transmission of the HRC VNs in all batches and supports decoding at the highest rate, which is infeasible in a Raptor code. Second, similar to the LT component in Raptor codes, XOR operations in the ESRL code generate the VNs of the IRC and the SSC, producing additional parity bits row-wise. However, these connections in the ESRL code are pre-defined rather than arbitrary. Third, compared to two-phase BP decoding of Raptor codes or joint BP decoding on PBRL codes, the ESRL code can support a more flexible windowed BP decoding to slide over the coupled protograph $\mathcal{G}_{\text{SC}}$ in a low-complexity manner [47]. More details on the decoding are discussed in Section V-C and Section VI-C.
@@ III-B Properties of ESRL Codes
In this section, we present some important properties of ESRL codes. First, the ESRL code is a time-invariant SC-LDPC code with no parallel edges, i.e., each entry of $\mathbf{B}$ is binary, and no non-zero entries in $\mathbf{B}_{i},\forall i\in[0,\omega]$ occupy the same position. The time-invariant nature and the absence of parallel edges ensure that the ESRL code can be decomposed into multiple independent identical engines that decode on different $\mathbf{U}$ blocks in parallel, without causing memory access conflicts (assuming the decoding schedule is consistent at each engine). This results in ESRL codes that allow for significant decoding parallelism to potentially achieve the high throughput demand of next-generation channel coding, as demonstrated in Section V-C and Section VI. It is noteworthy that although parallel edges can reduce the decoding threshold [18, 48] and can be easily assigned to $\omega$ sub-matrices in SC-LDPC codes without the need for two-step lifting [48], this performance gain sacrifices parallelism in an ESRL decoder. The absence of parallel edges, as mentioned in [38], can also simplify the theoretical tools described in Section IV to optimize ESRL codes.
The HRC protograph is the protograph design of a good code at a given high rate, which has been well studied in [48]. Following the example of [48], we set the minimum column weight to at least $3$ to ensure linear minimum distance growth with increasing blocklength. Besides, a punctured VN that is fully connected to all CNs in the HRC is placed in the first column to reduce the decoding threshold. The IRC component aims to provide reliability back to the VNs of the original HRC. In a complete ESRL matrix with the lowest rate, the CNs of the IRC eventually connect back to all VNs of the HRC. Following [49], each row of the IRC has at least two edges to retain linear minimum distance growth with increasing blocklength. Besides, the punctured VN of the precode connects to all CNs in the IRC. Unlike the use of parallel edges in [18], the HRC and the IRC of ESRL codes do not involve any parallel edges.
The SSC protograph is to enhance message propagation on $\mathbf{U}$, especially for medium to low rates. In the SSC, all diagonal elements are set to ‘$1$’s to facilitate encoding. Note that the SSC is connected to parity bits that are generally less critical than the VNs in the HRC, thus the number of edges in the SSC should be kept lower than that of the IRC. While the rate reduces, we increase the SSC weight to avoid the right side being too sparse, but by experimental trial, this proportion is at most $30\%$ of the IRC weight.
@@ III-C Code Profile of ESRL Codes
The uncoupled protomatrix $\mathbf{B}$ outlines the parent structure of ESRL codes, as explained in Section III-A. To convert $\mathbf{B}$ to $\mathbf{B}_{\text{SC}}$, we adopt an edge-spreading matrix $\mathbf{T}$ that was proposed in [31] and has been adopted in [34, 38], instead of explicitly recording $\omega+1$ sub-matrices as in [28, 29]. The matrix $\mathbf{T}$ of size $m\times n$ determines how the edges of the uncoupled protograph $\mathcal{G}$ are spread to the coupled protograph $\mathcal{G}_{\text{SC}}$. Specifically, the elements of $\mathbf{T}$ indicate to which sub-matrices the edges of $\mathbf{B}$ are assigned. We set $\mathbf{T}_{c,v}=-1$ if $\mathbf{B}_{c,v}=0$. Otherwise, $\mathbf{T}_{c,v}=i$, $i\in[0,\omega]$, implies that the entry at row $c$ and column $v$ of $\mathbf{B}_{c,v}$ is assigned to the corresponding position in $\mathbf{B}_{i}$. Let $\delta_{i}(\mathbf{T}_{c,v})$ be a function that returns a boolean value of ‘$1$’ if $\mathbf{T}_{c,v}$ equals $i$, and ‘00’ otherwise. As a result, we can generate $\omega+1$ sub-matrices $\mathbf{B}_{0},\mathbf{B}_{1},\dots,\mathbf{B}_{\omega}$ explicitly as
@@
Furthermore, the termination of finite-length SC-LDPC codes results in an unavoidable rate loss at the boundary. To mitigate this issue, one can either append a redundancy matrix $\mathbf{Q}$ at the end [50, 46] or reduce the information length of the last block [51] to satisfy encoding. ESRL codes adopt the former approach, as it has a lower rate loss. Previous works generally use a simple identity matrix for $\mathbf{Q}$ [50, 46], but this leads to serious error floors at medium to low rates. In contrast, we adopt a tri-diagonal matrix for $\mathbf{Q}$ with a lower triangular structure to ensure that most column degrees are either $2$ or $3$. For rate compatibility, the matrix $\mathbf{Q}$ must remain full-rank even after pruning. Based on the three matrices $\mathbf{B}$, $\mathbf{T}$, and $\mathbf{Q}$, we explicitly define the coupled protomatrix $\mathbf{B}_{\text{SC}}$ of the ESRL code as follows
@@ IV-A Unified Graph for ESRL Codes
Inspired by the aforementioned code profile, we propose a novel graph representation for ESRL codes, which we call the unified graph $\mathcal{G}_{\text{U}}$. This graph contains all essential code information and enables a comprehensive illustration of possible graph structures within an SC-LDPC protograph. Unlike traditional approaches, a single unified graph $\mathcal{G}_{\text{U}}$ can integrate the relationships and dependencies among the matrices $\mathbf{B}$ and $\mathbf{T}$ in ESRL codes, [^1] Given that the tail matrix $\mathbf{Q}$ has a fixed structure and is not involved in edge spreading, we ignore the presence of $\mathbf{Q}$ to simplify the optimization on $\mathcal{G}_{\text{U}}$. Note that $\mathcal{G}_{\text{U}}$ is a protograph concept. thereby offering a more global view of the code construction. The construction of $\mathcal{G}_{\text{U}}$ is built on the bipartite graph derived from $\mathbf{B}$, where edges are annotated by different line types based on the corresponding elements of $\mathbf{T}$. For example, Fig. 4 depicts the unified graph $\mathcal{G}_{\text{U}}$ derived from $\mathbf{B}$ and $\mathbf{T}$ with $\omega=1$, where dashed and solid lines distinguish whether edge $e$ is assigned to $\mathbf{B}_{0}$ or $\mathbf{B}_{1}$, respectively. In this section, we present a message-passing algorithm on $\mathcal{G}_{\text{U}}$ to count and eliminate short cycles of $\mathcal{G}_{\text{SC}}$ efficiently, which inspires a good spreading strategy on $\mathcal{G}_{\text{U}}$ to enhance the performance of ESRL codes.
@@ IV-A1 Cycle Counting over the Unified Graph
For rigor, we only consider simple cycles in this paper, i.e., tailless backtrackless closed (TBC) walks in a graph. After edge spreading from $\mathcal{G}$ to $\mathcal{G}_{\text{SC}}$, the number of $l$-cycles in $\mathcal{G}$ is the upper bound of the number of $l$-cycles in $\mathcal{G}_{\text{SC}}$ if repeated cycles in $L$ replicas are counted only once [38].
@@
Subsequently, we introduce a message-passing algorithm over $\mathcal{G}_{\text{U}}$ to efficiently count cycles for ESRL codes. In general, a message-passing algorithm computes messages at the nodes and propagates them along the edges of a graph, where the calculation of the message sent along an edge $e$ excludes the message previously received along $e$ (refers to this property as extrinsic[7]. Let $\gamma$ denote a basic message, and an example of an extrinsic message-passing process is illustrated in Fig. 4. Hence, the messages over the edges of $\mathcal{G}_{\text{U}}$ are all monomials of $\gamma$ and the message-passing operation is multiplication. Besides, we need to record the extra accumulated path value $\phi$ of each $\gamma$ over $\mathcal{G}_{\text{U}}$, due to using edge spreading. Let $t$ denote the index of discrete time, and we call each $t$ as one step, corresponding to either a VN-to-CN or CN-to-VN phase, where $t$ always starts from 00 of the VN-to-CN phase. We denote $\mathcal{M}_{v,c}^{(t)}$ as the VN-to-CN message along $e_{v}^{c}$ and $\mathcal{M}_{c,v}^{(t)}$ as the CN-to-VN message along $e_{c}^{v}$. The update rules are defined in (10) and (11), and the $\phi$ of each message is also tracked dynamically to verify whether the cycle remains or not in $\mathcal{G}_{\text{SC}}$ based on (9).
Herein, we present how to use $\mathcal{G}_{\text{U}}$ count cycles that include a specific VN. For the $v$-th VN, the initial message $\mathcal{M}_{v,c}^{(0)}$ at $t=0$ along the edge $e_{v}^{c}$ is denoted by $\gamma_{c}^{\phi_{c}}$, where $\phi_{c}$ is initialized as $\mathbf{T}_{c,v}$. We can summarize the VN-to-CN message-passing rules as
@@
Notably, the general message format $\mathcal{M}$ that propagates along the edges in $\mathcal{G}_{\text{U}}$ is the product of multiple $\gamma$ variables, each labeled by individual path value $\phi$, where both the number and type of $\gamma$ are dynamic. At the $t\!=\!l-1$-th step, we will obtain the number of $l$-cycles that include the $v$-th VN as
$Q_{v}^{l}=f\left(p_{v}^{l})\right/2,$ (14)
where the division by $2$ arises from the directed edges, as each cycle is counted twice. The function $f\left(p_{v}^{l}\right)$ computes the number of all $\gamma$ variables with $\phi=0$ in the polynomial $p_{v}^{l}=\sum_{c\in\mathbb{C}_{v}}\mathcal{M}_{c,v}^{(l-1)}$, i.e., only sequences with $\phi=0$ will remain as cycles in $\mathcal{G}_{\text{SC}}$ after edge spreading. Note that due to the extrinsic property mentioned before, the component formed by the initial message $\gamma_{c}$ sent along the edge $e_{v}^{c}$ should be excluded from $\mathcal{M}_{c,v}^{(l-1)}$ in (14[52]. Consequently, we repeat the above procedures for all VNs. The total number of $l$-cycles in the entire $\mathcal{G}_{\text{SC}}$ can be obtained as
$Q^{l}=\left(\sum\limits_{v\in\mathbb{V}}Q_{v}^{l}\right)/\left(l/2\right),$ (15)
where the accumulated result in (15) needs to be divided by $l/2$ as every $l$-cycle is counted $l/2$-times. Define the shortest cycle length in a graph as the girth $g$. The above approach can efficiently count $l$-cycles when $l\in\{4,6,\dots,2g-2\}$.
@@
    At $t=0$, $v_{4}$ sends the initial messages $\gamma_{1}^{1}$$\gamma_{2}^{0}$, and $\gamma_{3}^{0}$ along the adjacent edges $e_{v_{4}}^{c_{1}}$$e_{v_{4}}^{c_{2}}$, and $e_{v_{4}}^{c_{3}}$, respectively, where the initial path values are according to the corresponding elements of $\mathbf{T}$ as $\{1,0,0\}$. All the messages sent by $v_{0}$, $v_{1}$, $v_{2}$, and $v_{3}$ along their edges are equal to $1$ and omitted. At $t=1$, all valid messages are updated based on (11) and (13) for the CN-to-VN phase.